chemical eq module page 33 q1  and page 28 q1

Grade:11

1 Answers

Naveen Kumar
9 years ago
Q1, Page33
The given two equations are atequilibrium.
Then after adding Ni +CO......>Ni(CO)4
so Co will decease and to prevent the decrease(a/c to Le-Chat. Principle) COCl2 will dissociate and more Cl2 will be realeased.
This increase in Cl2 will push the first reaction backward and so PCl3 will decrease.

Q1, page 28
let partial pressure of NH3=P1=partial pressure ofH2S
and p1+p1=X(given)
Now kp=p12=(X/2)2
so free energy change=-RTlnk...........
................................=-RTln(X/2)2
.............................................
=-2RT(lnX-ln2)